OpenStudy (anonymous):

How do i plot y+3=x? please explain How do i plot y+3=x? please explain @Mathematics

OpenStudy (anonymous):

subtract 3 from both sides and you will get: y=x-3. Now it is in y=mx+b from, where m is the slope

OpenStudy (anonymous):

i dont understand???

OpenStudy (anonymous):

Now that you have y=x-3. You can find the x and y intercepts